OpenStudy (anonymous):

x(x-4)=-29 solve with quadratic equation

OpenStudy (anonymous):

Look : We have : x*x+x*(-4)=-29 K ?:)

OpenStudy (azureilai):

Distribute x, and balance equation \[x ^{2}-4x+29=0\] quadratic formula: \[\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a}\]

OpenStudy (anonymous):

@jacksond96